# Extended mean field control problems with constraints: The generalized Fritz-John conditions and Lagrangian method

### v1

@@ Abstract
This paper studies the extended mean field control problems under dynamic expectation constraints and/or dynamic state-control constraints. We aim to pioneer the establishment of the stochastic maximum principle (SMP) and the derivation of the backward SDE (BSDE) from the perspective of the constrained optimization using the method of Lagrangian multipliers. To this end, we first propose to embed the constrained extended mean-field control (C-(E)MFC) problems into some abstract optimization problems with constraints on Banach spaces, for which we develop the generalized Fritz-John (FJ) optimality conditions. We then prove the stochastic maximum principle (SMP) for C-(E)MFC problems by transforming the FJ type conditions into an equivalent stochastic first-order condition associated with a general type of constrained forward-backward SDEs (FBSDEs). Contrary to the existing literature, we recast the controlled Mckean-Vlasov SDE as an infinite-dimensional equality constraint such that the BSDE induced by the FJ first-order optimality condition can be interpreted as the generalized Lagrange multiplier to cope with the SDE constraint. Finally, we also present the SMP for constrained stochastic control problems and constrained mean field game problems as consequences of our main result for C-(E)MFC problems.
Keywords: Extended mean-field control, dynamic expectation constraints, dynamic state-control constraints, stochastic maximum principle, generalized Fritz-John optimality condition.
@@ 1 Introduction
Motivated by wide applications in finance, engineering and management science, stochastic control problems with various types of constraints have attracted tremendous attention over the past decades. Several types of constraints on state process and the control process have been predominantly investigated in different context such as the dynamic (in time) state and/or control constraint, the dynamic (in time) expectation constraint on state and/or control paths, and their simplified versions as static constraints only at the terminal time. New techniques are inevitably needed in order to address the constraints such that the dynamic programming approach and Pontryagin’s maximum principle approach can be modified and adopted.
To name a few in the direction of studies to cope with the state and/or control constraints, Hu and Zhou [27] investigate a stochastic LQ control problem with random coefficients by employing Tanaka’s formula, where the control variable is constrained in a cone; Bonnans and Silva [10] examine the stochastic control problems with control constraints and several equality and inequality constraints over the final state, for which the first order and second order optimality conditions are established; Liu et al. [31] consider stochastic control with an end-point constraint by establishing a general FJ condition in Banach space where only the control variable is reformulated as an optimization problem; Hu et al. [25] investigate a stochastic LQ control problem with regime switching, random coefficients and cone control constraint, where the optimal feedback control and optimal cost value are characterized explicitly via two systems of extended stochastic Riccati equations; Zhang and Zhang [42] study a stochastic linear-quadratic (LQ) control problem with a static equality constraint on the terminal state by using Lagrangian duality method.
@@
MFC problems, also called the McKean-Vlasov optimal control problems, feature the cooperative interactions in the mean field model where all agents jointly optimize the social optimum that yields the optimal control from the perspective of the social planner, which have raised a lot of research interests in the past few years. To name a few, Acciaio et al. [1] explore suitable versions of the Pontryagin stochastic maximum principle (SMP) for extended MFC problems in which the cost function and the state process depend on the joint distribution of the state and the control process; The first-order variation in Bensoussan [5] and the spike-variation method in Peng [35] have been generalized to cope with MFC problems among [2, 7, 33, 21] that handle McKean-Valsoc dynamics with coefficients depending on the moments of the population state distribution; Li [30] examines MFC problems and establish the SMP in scalar interaction forms; Bensoussan et al. [6] introduce a unified approach for MFC and MFG problems by using the HJB-FP coupled equations and the SMP; Carmona and Delarue [14] solve a class of MFC problems with a general dependence on the law of the state process; Pham and Wei [38] investigate the dynamic programming principle for MFC problems in the presence of common noise; Pham and Wei [39] examine the extended MFC with state-control joint law dependence by reformulating the problem into a deterministic control problem with only the marginal distribution of the process as controlled state variable; Bo et al. [8] consider a LQ-MFC problem by proving the Gamma convergence of the optimal controls from the N-player games to the mean-field limit model; Nie and Yan [34] study extended MFC problems under partial information; Bo et al. [9] recently establish the SMP and HJB eqaution for an extended MFC problem with state-control joint law and Poissonian common noise by proposing the extension transformation.
@@
MFC and MFG problems under the dynamic expectation constraints and/or the dynamic state-control constraints are relatively under developed. Grammatico et al. [24] study a deterministic MFC problem in the presence of heterogeneous convex constraints at the level of individual agents; Du and Wu [22] consider a LQG MFC problem with control constraints; Bonnans et al. [12] analyze the existence of equilibria for a class of deterministic MFGs of controls with state and control constraints at the terminal time by using Kahutani’s theorem; Daudin [19] study the convergence problem of mean-field control theory in the presence of state constraints; Daudin [20] examine a problem of optimal control of the Fokker-Planck equation with state constraints in the Wasserstein space and give first-order necessary conditions for optimality in the form of a mean-field game system of partial differential equations associated with an exclusion condition; Craig et al. [18] discuss a class of first-order MFC problem with source and terminal constraints by using a particle method; Meherrem and Hafayed [32] prove Peng’s necessary optimality conditions for a general mean-field system under state constraints based on conventional variational principle.
@@
In the present paper, we study extended mean field control problems under both the dynamic expectation constraints and the dynamic state-control constraints from the perspective of optimization problems with Banach space valued constraints. We propose a new methodology that bridges the SMP method for the C-(E)MFC problems and the Lagrange multipliers method with the generalized Fritz-John (FJ) optimality condition. In a nutshell, when the forward McKean-Vlasov state process is taken as an infinite-dimensional constraints, the generalized FJ optimality condition for the control variable yields the stochastic (first-order) minimum condition that leads to the SMP. Meanwhile, the generalized Lagrangian multiplier in the space of Itô processes induced by the FJ optimality condition on the state variable gives us the adjoint process as the solution to the BSDE.
@@
More concretely, we first develop the type of FJ optimality condition (Theorem 2.7) for nonlinear optimization problems with (Banach space valued) equality and/or inequality constraints by using the generalized gradient introduced in Clarke [17]. Next, as the main novelty of this paper, we formulate our C-(E)MFC problem and transform it into an abstract optimization problem with constraints on Banach spaces (see problem (26)), for which we can apply the newly developed FJ optimality condition on Banach spaces. We then contribute to the establishment of SMP for the C-(E)MFC problem (Theorem 3.11) by transforming the FJ conditions into a stochastic first-order condition that gives rise to a generalized type of constrained FBSDEs. In sharp contrast to the conventional methods, we broaden the scope by considering both state and control as targeted variables in the formulation of the constrained optimization problems such that the Lagrangian multipliers method can be employed. Furthermore, we propose to treat the controlled Mckean-Vlasov state dynamics as an (integral) equality constraint on an infinite dimensional (Hilbert) space, which is also new to the literature. Then, the adjoint process as a solution to a generalized type of constrained BSDEs induced by the FJ first-order optimality condition can be interpreted as the generalized Lagrangian multiplier to handle the forward SDE equality constraint, for which a new space denoted by $\mathbb{M}^{2}$ is introduced, which consists of all Itô processes to characterize the integration formulation of the Mckean-Vlasov type SDE constraint. As special cases of our C-(E)MFC problems, we also derive and present the SMP results for the stochastic optimal control problems as well as mean field game (MFG) problem with the same types of constraints.
@@
The rest of the paper is organized as follows. Section 2 gives an overview of the generalized gradients on Banach space in [17] and establishes the generalized FJ optimality condition and the constraint qualification (CQ) for a class of constrained optimization problems on Banach spaces. Section 3 reformulates our C-(E)MFC as an abstract constrained optimization problem on Banach spaces and develops the SMP and derives the general type of constrained FBSDEs by using the generalized FJ optimality condition. Section 4 presents the SMP results for stochastic optimal control problems and extended MFG problems with the same types of dynamic expectation constraints and state-control constraints as special cases of the main results from Section 3. Finally, the proofs of some auxiliary results in previous sections are reported in Appendix A.
@@
$D_{X}f$ The partial Fréchet derivative of $f$ with resepct to $X$
$D_{\alpha}f$ The partial Fréchet derivative of $f$ with resepct to $\alpha$
@@
$\mathbb{E}$ Expectation operator under probability measure $\mathbb{P}$
$\mathbb{E}^{\prime}$ Expectation operator under probability measure $\mathbb{P}^{\prime}$
@@ 2 Generalized Fritz-John Optimality Condition
For the goal above, we propose some techniques based on generalized gradients introduced in Clarke [17] in which the constraints are only restricted in Euclidean space in the sense that $\mathtt{Y}=\mathtt{H}=\mathbb{R}$ and $K_{i}=\mathbb{R}_{+}\cup\{0\}$ for all $i\in I$. The general Banach space valued constraints formulated in problem (1) are customized to study the C-(E)MFC problems that will be discussed in Section 3.
@@ 2.1 A review of generalized gradients on Banach space
We then can show the next result.
@@
Let $\mathcal{I}$ be an index set and $(\phi_{i})_{i\in\mathcal{I}}$ be a family of uniformly local Lipschitz (real-valued) functions on $\mathtt{X}$ (i.e., their Lipschitz constants are independent of $i\in\mathcal{I}$). Define the mapping $\phi:\mathtt{X}\mapsto\mathbb{R}$ by $\phi(x)=\sup_{i\in\mathcal{I}}\phi_{i}(x)$ for $x\in\mathtt{X}$, and assume that $M(x):=\{i\in\mathcal{I};~{}\phi_{i}(x)=\phi(x)\}\neq\varnothing$. Then, $x\mapsto\phi(x)$ is locally Lipschitz and it holds that
$\partial\phi(x)\subset\overline{\mathrm{co}\{\partial\phi_{i}(x);~{}i\in M(x)\}}.$
Proof.
The first claim is straightforward. It suffices to show the second assertion. Denote by $B$ the nonempty set on the right side of the second assertion. By the uniformly local Lipschitz continuity of $\phi_{i}:\mathtt{X}\mapsto\mathbb{R}$ with $i\in\mathcal{I}$, we conclude that $B$ is a closed, convex and bounded set, and thus it is convex and weak- compact. Furthermore, by Lemma 2.1-(ii), it holds that, for any $v\in\mathtt{X}$,
$\sup_{\xi\in B}\langle v,\xi\rangle_{\mathtt{X},\mathtt{X}^{*}}=\sup_{i\in M(x)}\phi_{i}^{\rm o}(x;v),$
where $\phi_{i}^{\rm o}(x;v)$ is the generalized directional derivative of $\phi_{i}$ in the direction $v$ given in (2), we can directly show from the definition that $\phi^{\rm o}(x;v)$ is no larger than the right-hand side of the above equality. Then, the desired result follows from Lemma 2.1-(iii). ∎
@@
We have the following dual relationship between $K_{i}\subset\mathtt{Y}_{i}$ and $K_{i}^{+}\subset\mathtt{Y}_{i}^{*}$.
@@
Proof.
For $i\in I$, denote $L_{i}$ as the set on the right hand side of (5). Then, it follows from the definition of $K_{i}^{+}$ that $K_{i}\subset L_{i}$. We will argue by contradiction to show $L_{i}\subset K_{i}$. In fact, if there exists a point $y\in L_{i}$ and $y\notin K_{i}$, the separation theorem yields that, there exists a $\xi\in\mathtt{Y}_{i}^{*}$ such that
$\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}<0\leq\langle x,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}},\quad\forall x\in K_{i}.$
This leads to $\xi\in K_{i}^{+}$, and hence $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\geq 0$ due to the definition of $L_{i}$, which yields a contradiction. ∎
We also recall two important properties on Banach spaces as below.
Lemma 2.5.
For a Banach space $(\mathtt{X},\|\cdot\|_{\mathtt{X}})$, the following results hold:
  • (i)
    If $\mathtt{X}$ is separable, any bounded set in $\mathtt{X}^{*}$ is weak- sequentially compact;
  • (ii)
    If $\mathtt{X}$ is reflexive, any bounded set in $\mathtt{X}^{*}$ is weak- sequentially compact.
@@
Next, we argue by contradiction to show that the other side of the claim holds. Suppose that $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}>0$ for all nonzero $\xi\in K_{i}^{+}$ and $y\notin K_{i}^{\rm o}$, which implies $y\in\overline{K_{i}^{c}}$ with $K_{i}^{c}$ being the complement of $K_{i}$. Consequently, there exists a sequence $(y_{k})_{k\in\mathbb{N}}\subset K_{i}^{c}$ such that $y_{k}\to y$ as $k\to\infty$. Then, in light of Lemma 2.4, there exists a $\xi_{k}\in K_{i}^{+}$ such that $\langle y_{k},\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}<0$ for all $k\in\mathbb{N}$. We assume without loss of generality that $\|\xi_{k}\|_{\mathtt{Y}_{i}^{*}}=1$. Thus, by Lemma 2.5-(i), there exists a weak- convergent subsequence of $(\xi_{k})_{k\in\mathbb{N}}$ (still denoted by $(\xi_{k})_{k\in\mathbb{N}}$) such that $\xi_{k}\to\xi\in\mathtt{Y}_{i}^{*}$ in the topology of weak- convergence and $\xi\in K_{i}^{+}$ results from the definition of $K_{i}^{+}$ Hence, it holds that
$\left|\langle y_{k},\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}-\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|$ $\leq\left|\langle y_{k}-y,\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|+\left|\langle y,\xi_{k}-\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|$
$\leq\|y_{k}-y\|_{\mathtt{Y}_{i}}+\left|\langle y,\xi_{k}-\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|.$

### v2

@@ Abstract
This paper studies the extended mean field control problems under general dynamic expectation constraints and/or dynamic pathwise state-control and law constraints. We aim to pioneer the establishment of the stochastic maximum principle (SMP) and the derivation of the backward SDE (BSDE) from the perspective of the constrained optimization using the method of Lagrangian multipliers. To this end, we first propose to embed the constrained extended mean-field control (C-MFC) problems into some abstract optimization problems with constraints on Banach spaces, for which we develop the generalized Fritz-John (FJ) optimality conditions. We then prove the stochastic maximum principle (SMP) for C-MFC problems by transforming the FJ type conditions into an equivalent stochastic first-order condition associated with a general type of constrained forward-backward SDEs (FBSDEs). Contrary to the existing literature, we treat the controlled Mckean-Vlasov SDE as an infinite-dimensional equality constraint such that the BSDE induced by the FJ first-order optimality condition can be interpreted as the generalized Lagrange multiplier to cope with the SDE constraint. Finally, we also present the SMP for stochastic control problems and mean field game problems under similar types of constraints as consequences of our main result for C-MFC problems.
Keywords: Extended mean-field control, dynamic expectation constraints, dynamic state-control and law constraints, stochastic maximum principle, generalized Fritz-John optimality condition.
2020 MSC: 93E20, 49N80, 46N10, 60H10
@@ 1 Introduction
Motivated by wide applications in finance, engineering and management science, stochastic control problems with various types of constraints have attracted tremendous attention over the past decades. Several types of constraints on state process, the control process and their laws have been predominantly investigated in different context such as the dynamic pathwise state, control and law constraints, the dynamic expectation constraints on state and/or control paths, and their simplified versions as static constraints only at the terminal time. New techniques are inevitably needed to address these constraints in order to employ the dynamic programming approach or the Pontryagin maximum principle approach.
To name a few in the direction of studies to cope with the pathwise state, control and law constraints, Hu and Zhou [28] investigate a stochastic LQ control problem with random coefficients by employing Tanaka’s formula, where the control variable is constrained in a cone; Bonnans and Silva [10] examine the stochastic control problems with control constraints and several equality and inequality constraints over the final state, for which the first order and second order optimality conditions are established; Liu et al. [32] consider stochastic control with an end-point constraint by establishing a general FJ condition in Banach space where only the control variable is reformulated as an optimization problem; Hu et al. [26] investigate a stochastic LQ control problem with regime switching, random coefficients and cone control constraint, where the optimal feedback control and optimal cost value are characterized explicitly via two systems of extended stochastic Riccati equations; Zhang and Zhang [43] study a stochastic linear-quadratic (LQ) control problem with a static equality constraint on the terminal state by using the Lagrangian duality method.
@@
MFC problems, also called the McKean-Vlasov optimal control problems, feature the cooperative interactions in the mean field model where all agents jointly optimize the social optimum that yields the optimal control from the perspective of the social planner, which have raised a lot of research interests in the past few years. To name a few, Acciaio et al. [1] explore suitable versions of the Pontryagin stochastic maximum principle (SMP) for extended MFC problems in which the cost function and the state process depend on the joint distribution of the state and the control process; The first-order variation in Bensoussan [5] and the spike-variation method in Peng [36] have been generalized to cope with MFC problems among [2, 7, 34, 21] that handle McKean-Valsoc dynamics with coefficients depending on the moments of the population state distribution; Li [31] examines MFC problems and establish the SMP in scalar interaction forms; Bensoussan et al. [6] introduce a unified approach for MFC and MFG problems by using the HJB-FP coupled equations and the SMP; Carmona and Delarue [14] solve a class of MFC problems with a general dependence on the law of the state process; Pham and Wei [39] investigate the dynamic programming principle for MFC problems in the presence of common noise; Pham and Wei [40] examine the extended MFC with state-control joint law dependence by reformulating the problem into a deterministic control problem with only the marginal distribution of the process as controlled state variable; Bo et al. [8] consider a LQ-MFC problem by proving the Gamma convergence of the optimal controls from the N-player games to the mean-field limit model; Nie and Yan [35] study extended MFC problems under partial information; Bo et al. [9] recently establish the SMP and HJB eqaution for an extended MFC problem with Poissonian common noise by proposing the extension transformation.
@@
MFC and MFG problems under the dynamic expectation constraints and/or the dynamic state-control and law constraints are relatively under developed. Grammatico et al. [25] study a deterministic MFC problem in the presence of heterogeneous convex constraints at the level of individual agents; Du and Wu [22] consider a LQG MFC problem with control constraints; Germain et al. [24] apply the level-set approach to some MFC problems under dynamic state-law constraints and discuss some applications in renewable energy storage and portfolio selections; Bonnans et al. [12] analyze the existence of equilibria for a class of deterministic MFGs of controls with state and control constraints at the terminal time by using Kahutani’s theorem; Daudin [19] study the convergence problem of mean-field control theory in the presence of state constraints; Daudin [20] examine a problem of optimal control of the Fokker-Planck equation with state constraints in the Wasserstein space and give first-order necessary conditions for optimality in the form of a mean-field game system of partial differential equations associated with an exclusion condition; Craig et al. [18] discuss a class of first-order MFC problem with source and terminal constraints by using a particle method; Meherrem and Hafayed [33] prove Peng’s necessary optimality conditions for a general mean-field system under state constraints based on conventional variational principle.
@@
In the present paper, we study extended mean field control problems under both the dynamic expectation constraints and the pathwise state-control and law constraints from the perspective of optimization problems with Banach space valued constraints. We propose a new methodology that bridges the SMP method for the C-MFC problems and the Lagrange multipliers method based on the generalized Fritz-John (FJ) optimality condition. In a nutshell, when the forward McKean-Vlasov state process is taken as an infinite-dimensional constraints, the generalized FJ optimality condition for the control variable yields the stochastic (first-order) minimum condition that leads to the SMP. Meanwhile, the generalized Lagrangian multiplier in the space of Itô processes induced by the FJ optimality condition on the state variable gives us the adjoint process as the solution to the BSDE.
@@
More concretely, we first develop the type of FJ optimality condition (Theorem 2.6) for nonlinear optimization problems with (Banach space valued) equality and/or inequality constraints by using the generalized gradient introduced in Clarke [17]. Next, as the main novelty of this paper, we formulate our C-MFC problem and transform it into an abstract optimization problem with constraints on Banach spaces (see problem (19)), for which we can apply the newly developed FJ optimality condition on Banach spaces. We then contribute to the establishment of SMP for the C-MFC problem (Theorem 3.12) by transforming the FJ conditions into a stochastic first-order condition that gives rise to a generalized type of constrained FBSDEs. In sharp contrast to the conventional methods, we broaden the scope by considering both state and control as targeted variables in the formulation of the constrained optimization problems such that the Lagrangian multipliers method can be employed. Furthermore, we propose to treat the controlled Mckean-Vlasov state dynamics as an (integral) equality constraint on an infinite dimensional Hilbert space, which is also new to the literature. Then, the adjoint process as a solution to a generalized type of constrained BSDEs induced by the FJ first-order optimality condition can be interpreted as the generalized Lagrangian multiplier to handle the forward SDE equality constraint, for which a new space denoted by $\mathbb{M}^{2}$ is introduced, which consists of all Itô processes to characterize the integration formulation of the Mckean-Vlasov type SDE constraint. As special cases of our C-MFC problems, we also present the SMP results for the stochastic optimal control problems as well as mean field game (MFG) problem with the same types of dynamic constraints.
@@
The rest of the paper is organized as follows. Section 2 gives an overview of the generalized gradients on Banach space and establishes the generalized FJ optimality condition and the constraint qualification (CQ) for a class of constrained optimization problems on Banach spaces. Section 3 reformulates our C-MFC as an abstract constrained optimization problem on Banach spaces and develops the SMP and the associated constrained FBSDEs by using the generalized FJ optimality condition. Section 4 presents the SMP results for constrained stochastic optimal control problems and constrained MFG problems as special cases of the main results from Section 3. Finally, the proofs of some auxiliary results in are reported in Appendix A.
@@
$D_{X}f$ ($D_{\alpha}f$) The partial Fréchet derivative of $f$ with resepct to $X$ ($\alpha$)
@@
$\mathbb{E}$ ($\mathbb{E}^{\prime}$) Expectation operator under probability measure $\mathbb{P}$ ($\mathbb{P}^{\prime}$)
@@ 2 Generalized Fritz-John Optimality Condition
For the goal above, we propose some techniques based on generalized gradients introduced in Clarke [17] in which the constraints are only restricted in Euclidean space in the sense that $\mathtt{Y}=\mathtt{H}=\mathbb{R}$ and $K_{i}=\mathbb{R}_{+}\cup\{0\}$ for all $i\in I$. The general Banach space valued constraints formulated in problem (1) are customized to study the C-MFC problems that will be discussed in Section 3.
@@ 2.1 A review of generalized gradients on Banach space
Using Lemma 2.1, we can show the next result.
@@
Let $\mathcal{I}$ be an index set and $(\phi_{i})_{i\in\mathcal{I}}$ be a family of uniformly local Lipschitz (real-valued) functions on $\mathtt{X}$ (i.e., their Lipschitz constants are independent of $i\in\mathcal{I}$). Define the mapping $\phi:\mathtt{X}\mapsto\mathbb{R}$ by $\phi(x)=\sup_{i\in\mathcal{I}}\phi_{i}(x)$ for $x\in\mathtt{X}$, and assume that $M(x):=\{i\in\mathcal{I};~{}\phi_{i}(x)=\phi(x)\}\neq\varnothing$. Then, $x\mapsto\phi(x)$ is locally Lipschitz and it holds that $\partial\phi(x)\subset\overline{\mathrm{co}\{\partial\phi_{i}(x);~{}i\in M(x)\}}.$
@@
Then, it is not difficult to show the following dual relationship between $K_{i}\subset\mathtt{Y}_{i}$ and $K_{i}^{+}\subset\mathtt{Y}_{i}^{*}$:
@@
Next, we argue by contradiction to show that the other side of the claim holds. Suppose that $\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}>0$ for all nonzero $\xi\in K_{i}^{+}$ and $y\notin K_{i}^{\rm o}$, which implies $y\in\overline{K_{i}^{c}}$ with $K_{i}^{c}$ being the complement of $K_{i}$. Consequently, there exists a sequence $(y_{k})_{k\in\mathbb{N}}\subset K_{i}^{c}$ such that $y_{k}\to y$ as $k\to\infty$. Then, in light of Lemma 2.4, there exists a $\xi_{k}\in K_{i}^{+}$ such that $\langle y_{k},\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}<0$ for all $k\in\mathbb{N}$. We assume without loss of generality that $\|\xi_{k}\|_{\mathtt{Y}_{i}^{*}}=1$. Thus, by the separability, there is a weak- convergent subsequence of $(\xi_{k})_{k\in\mathbb{N}}$ (still denoted by $(\xi_{k})_{k\in\mathbb{N}}$) such that $\xi_{k}\to\xi\in\mathtt{Y}_{i}^{*}$ in the topology of weak- convergence and $\xi\in K_{i}^{+}$ results from the definition of $K_{i}^{+}$ Hence
$\left|\langle y_{k},\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}-\langle y,\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|$ $\leq\left|\langle y_{k}-y,\xi_{k}\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|+\left|\langle y,\xi_{k}-\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|\leq\|y_{k}-y\|_{\mathtt{Y}_{i}}+\left|\langle y,\xi_{k}-\xi\rangle_{\mathtt{Y}_{i},\mathtt{Y}_{i}^{*}}\right|.$